OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (4, -4) with 0\(^o\) \(\le\) \(\theta\) < 360\(^o\).

Parth (parthkohli):

it's clear that \(r = 4\sqrt 2\). let's look at the angle.

Parth (parthkohli):

the angle here is \(\theta = 315^{\circ}\) (fourth quadrant).

Parth (parthkohli):

not sure what this thing means by two pairs, but i'd assume that the two pairs are a result of inverse tan.

Parth (parthkohli):

\[y/x = \tan \theta\]

OpenStudy (freckles):

think 315-180 puts us in the second quadrant  think about this other point in the second quadrant I drew